# Provably Efficient RLHF Pipeline: A Unified View from Contextual Bandits

### v1

@@ 4.2 Active Data Collection
$(x_{t},a_{t},a_{t}^{\prime})=\operatorname*{arg\,max}_{x,a,a^{\prime}\in\mathcal{X}\times\mathcal{A}\times\mathcal{A}}\left\{\big{\lVert}\phi(x,a)-\phi(x,a^{\prime})\big{\rVert}_{\mathcal{H}_{t}^{-1}}\right\}.$ (13)
Intuitively, $\mathcal{H}_{t}$ captures the directions explored so far, the selection criterion essentially seeks the tuple whose feature difference is poorly covered by the past observations.
Policy Output.
@@
After $T$ rounds, we define the reward function as $\widetilde{r}_{T}(x,a)=\frac{1}{T}\sum_{t=1}^{T}\phi(x,a)^{\top}\widetilde{\theta}_{t}$ as the average of all the past estimations. Then, the final policy is given by
$\pi_{T}(x)=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{r}_{T}(x,a).$ (14)
@@
Let $\delta\in(0,1]$, set the parameters as in Lemma 1, with probability at least $1-\delta$, Algorithm 2 ensures the following sub-optimality gap:
$\operatorname*{\mathsf{SubOpt}}(\pi_{T})\leq\widetilde{\mathcal{O}}\big{(}d\sqrt{{\kappa}/{T}}\big{)}.$
@@
In contrast to the passive data collection setting, the sub-optimality gap in Theorem 2 does not depend on the concentratability coefficient term. Instead, it decays at a rate of $T^{-1/2}$. This improvement arises because the active setting enables our algorithm to actively select queries with the maximum uncertainty, ensuring the uncertainty across all directions is effectively controlled.
@@
It is important to note that our result has a dependence on $\sqrt{\kappa}$. Though recent advances in the logistic bandits (Abeille et al., 2021; Zhang and Sugiyama, 2023) have shown that this dependence can be eliminated in the dominant term and shifted to lower-order terms, we believe this term remains necessary in our setting. This is due to the fact that the sub-optimality gap in our setting is defined as the difference between the reward $z^{\top}\theta^{*}$ while they focus the probability difference, i.e., $\sigma(z^{\top}\theta^{*})$. This difference prevents us from removing the $\sqrt{\kappa}$ term in our setting.
Comparison with Das et al. (2024).
Our algorithm attains the same sub-optimality gap as the work of Saha (2021), but with a significantly improved computational efficiency. Specifically, our algorithm has a time complexity of $\mathcal{O}(T)$ for the total $T$ round, which is more efficient than the $\mathcal{O}(T^{2}\log T)$ complexity of their MLE estimator. Moreover, our algorithm does not require storing the historical samples, resulting in only an $\mathcal{O}(1)$ storage complexity. In contrast, their MLE estimator necessitates storing all historical samples, leading to an $\mathcal{O}(T)$ storage complexity.
5 Deployment Stage
In the deployment stage, the learner uses the trained model to generate responses to user inputs and simultaneously gathering informative feedback to improve the model. This contrasts with the active data collection phase during training, which is focused solely on optimizing the final policy. In deployment, however, the learner is faced with a dual objective: selecting actions that maximize rewards to ensure a positive user experience, while also choosing actions that provide informative feedback for further model improvement.
Performance Measure.
To this end, we consider the cumulative regret compared to the optimal policy, which is defined as
${\operatorname*{Reg}}_{T}=\sum_{t=1}^{T}\left(r\left(x_{t},\pi^{*}(x_{t})\right)-\frac{r\left(x_{t},a_{t}\right)+r\left(x_{t},a^{\prime}_{t}\right)}{2}\right),$
where $\pi^{*}$ is the optimal policy, $a_{t}$ and $a^{\prime}_{t}$ are the actions selected by the learner at round $t$. This measure is stronger than the measure defined in Ji et al. (2024), which considers the gap between the optimal action and a single selected action. In the deployment phase, our proposed measure is more appropriate, as both actions must be sufficiently effective to ensure a positive user experience.
While our measure is more suitable for deployment, it also introduces significant challenges. To optimize the measure of Ji et al. (2024), the learner can exploit one action with high reward for exploitation while using the other action for exploration. In contrast, our setting requires both actions to be good enough. To address this, we propose a novel algorithm that balances exploration and exploitation for deployment. At a high level, our algorithm consists of two key components: estimator construction and query selection. We describe each component in detail below.
Estimator Construction.
The estimator construction is the same as the process used in the training stage. For each round $t\in[T]$, given $\left\{\left(x_{i},a_{i},a_{i}^{\prime},y_{i}\right)\right\}_{i=1}^{t-1}$, we construct the estimator $\widetilde{\theta}_{t}$ by OMD in an one-pass manner as in Eq. (9). Specifically, the estimator is updated by
$\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{2\eta}\big{\|}\theta-\widetilde{\theta}_{t}\big{\|}_{\widetilde{\mathcal{H}}_{t}}^{2}\Big{\}}.$ (15)
Query Selection.
Unlike in the training stage, during deployment, the learner must balance exploration and exploitation to minimize regret. Specifically, the learner must select queries that are both informative and likely to yield high rewards. Focusing solely on high-reward queries may cause the learner to miss valuable feedback for model improvement, while prioritizing only informative queries could prevent reward maximization. To address this, we propose a novel strategy inspired by Saha (2021), which accounts for both reward maximization and uncertainty maximization.
Based on the estimator $\theta_{t}$, we first define the promising set $\mathcal{A}_{t}$ that only contains the actions with potential to be the best action. Specifically, we define the promising set as
$\mathcal{A}_{t}=\big{\{}a\in\mathcal{A}\mid(\phi(x_{t},a)-\phi(x_{t},a^{\prime}))^{\top}\widetilde{\theta}_{t}+\widetilde{\beta}_{t}\lVert\phi(x_{t},a)-\phi(x_{t},a^{\prime})\rVert_{\mathcal{H}_{t}^{-1}}\geq 0,\forall a^{\prime}\in\mathcal{A}\big{\}}.$ (16)
Then, we select the action pairs with the maximum uncertainty in the promising set, i.e.,
$(a_{t},a_{t}^{\prime})=\operatorname*{arg\,max}_{(a,a^{\prime})\in\mathcal{A}_{t}\times\mathcal{A}_{t}}\left\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\right\|_{\mathcal{H}_{t}^{-1}}.$ (17)
Algorithm 3 Deployment Stage
0:  Regularization $\lambda>0$, failure probability $\delta\in(0,1]$
1:  Initialize $\widetilde{\theta}_{1}=0$, $\mathcal{H}_{1}=\lambda I_{d}$.
2:  for $t=1,2,\ldots,T$ do
3:     Observes context $x_{t}$.
4:     Compute the promising set by Eq. (16)
5:     Choose duel actions by Eq. (17).
6:     Observe preference feedback $y_{t}$
7:     Update ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
8:     Compute $\widetilde{\theta}_{t+1}$ by Eq. (15)
9:     Update $\mathcal{H}_{t+1}=\mathcal{H}_{t}+H_{t}(\widetilde{\theta}_{t+1})$
10:  end for
Theorem 3.
Let $\delta\in(0,1]$, set parameters as in Lemma 1, with probability at least $1-\delta$, Algorithm 3 enjoys the following regret guarantee:
Comparison with Saha (2021).
Our result improves upon theirs in both computational and statistical efficiency. For statistical efficiency, our algorithm achieves a regret of $\widetilde{\mathcal{O}}\big{(}d\sqrt{{\kappa}{T}}\big{)}$, which is tighter by a factor of $\sqrt{\kappa}$ compared to their $\widetilde{\mathcal{O}}\big{(}d{\kappa}\sqrt{{T}}\big{)}$ result. For computational efficiency, our algorithm updates in an online manner, with $\mathcal{O}(1)$ storage and $\mathcal{O}(T)$ time complexity for the total $T$ rounds. In contrast, their algorithm requires storing all the data and solving an optimization problem each round, leading to $\mathcal{O}(T)$ storage and $\mathcal{O}(T^{2}\log T)$ time complexity. As $T$ is large or even infinite during deployment, our algorithm has a significant advantage in computational efficiency.
6 Experiment
@@
In this section, we empirically evaluate the performance of our proposed method in terms of both effectiveness and efficiency. Specifically, we empirically validate our method with LLMs, and construct an online RLHF pipeline to evaluate different methods in both the training and deployment stage. [^1] Equal contribution. Next, we first describe the experimental setup, and then present the empirical results.
Experimental Setup.
@@
In our experiments, we employ the Llama-3-8B-Instruct  [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct as the foundation model for reward model. We extract features $\phi(x,a)$ using last layer of the model, and the feature space dimension is $d=4096$. We use the Ultrafeedback-binarized dataset, a pro-processed version of the original Ultrafeedback dataset (Cui et al., 2023) for evaluation, which is a large-scale, fine-grained, diverse preference dataset, widely used as a benchmark for RLHF. It collects about 64, 000 prompts from diverse resources, including question answering, summarization, and dialogue generation. Each data consists of a context $x$, two responses $a$ and $a^{\prime}$, and a human preference label $y$.
@@
11footnotetext: Our online RLHF pipeline code is available at https://github.com/ZinYY/RLHF_Pipeline
Implementation Details.
All experiments in this section were performed using four Nvidia A800 (80 GB) GPUs with two Intel(R) Xeon(R) Gold 6430 CPUs. The learning rate is set to $1\times 10^{-3}$ for all methods. We use DeepSpeed (Rasley et al., 2020) with mixed-precision training (BF16). The maximum sequence length is set to 1024 tokens, and we use a global batch size of 32. For practical efficiency, we implement our method using Hessian-vector product (HVP) and conjugate gradient methods, which efficiently approximate the Hessian matrix without the need for computing $d^{2}$ matrix. We use the default partition of the Ultrafeedback-binarized dataset with the standard train_prefs split for training and test_prefs split for evaluation. Furthermore, we also enable flash attention for the improved training efficiency.
6.1 Training Stage
@@
Figure 1: For the training stage with passive data collection, we compare our method with MLE (Zhu et al., 2023). We report the average accuracy and loss of the reward model during training.
In this part, we conduct experiments to evaluate the performance of our method in the training stage, including both passive and active data collection strategies.
Contenders.
In the training stage, we compare with:
  • Random: A baseline that randomly selects queries.
  • Passive-MLE (Zhu et al., 2023): A method that uses maximum likelihood estimation to learn the reward function and compute the pessimistic policy based on the learned reward model.
  • Active-MLE (Das et al., 2024): A method using MLE to optimize the reward model and selects queries based on uncertainty, i.e., the maximum difference of two actions.
We compared our method with other methods in terms of the accuracy of the reward model, as the method for training the final policy based on the learned reward model is the same as contenders.
Passive Data Collection Results.
We first evaluate our method in the passive data collection setting. We randomly sample $T=30,000$ data points from the Ultrafeedback dataset for training. Figure 1 shows the loss and accuracy vs. the number of training samples. We observe that our method converges faster to a lower loss and achieves a higher evaluation accuracy compared to baselines. The improvement is particularly pronounced in the small-sample regime ($T<10,000$), where our method achieves a higher evaluation accuracy with the same amount of samples compared to Passive-MLE which employs conventional stochastic gradient descent (SGD) updates. This demonstrates the superior statistical efficiency of our approach, therefore achieving a better performance with fewer training samples.
Active Data Collection Results.
We then evaluate our method in the active data collection setting. In this scenario, the global batch size is set to 8, and we only allow the method to select 6, 400 samples out of the whole training datasets for training according to different selection strategies. Table 2 and Figure 2 demonstrate that our active learning strategy achieves comparable performance to passive learning while using only 21% of the labeled data. Specifically, our method achieves an average accuracy of 70.43% after 6,400 actively selected samples, which is comparable with the previous state-of-the-art method, Active-MLE, while our method is 3 times faster because we do not need to recompute all the selected data in the buffer, but we update the model in an online manner. The training loss curves in Figure 2(a) also show that our method converges faster compared to other contenders. This improved sample efficiency can be attributed to our theoretically-motivated query selection criterion that effectively identifies the most informative samples, which further validates the two key factors in our method: (1) the use of local norm $\mathcal{H}_{t}$ that better captures the uncertainty in different directions, and (2) the computationally efficient one-pass update rule that enables processing large-scale datasets.

### v2

@@ 5.2 Online RLHF with Active Data Collection
$(x_{t+1},a_{t+1},a_{t+1}^{\prime})=\operatorname*{arg\,max}_{x,a,a^{\prime}\in\mathcal{X}\times\mathcal{A}\times\mathcal{A}}\left\{\big{\lVert}\phi(x,a)-\phi(x,a^{\prime})\big{\rVert}_{\mathcal{H}_{t+1}^{-1}}\right\}.$ (7)
@@
Policy Optimization.  After $T$ rounds, we define the reward as the average of all the past estimations $\widetilde{r}_{T+1}(x,a)=\frac{1}{T+1}\sum_{t=1}^{T+1}\phi(x,a)^{\top}\widetilde{\theta}_{t}$. The policy is given by $\pi_{T+1}(x)=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{r}_{T+1}(x,a)$.
@@
Set parameters as in Lemma 1, with probability at least $1-\delta$, Algorithm 3 ensures
$\operatorname*{\mathsf{SubOpt}}(\pi_{T+1})=\mathbb{E}_{x\sim\rho}\left[r(x,\pi^{*}(x))-r(x,\pi_{T+1}(x))\right]\leq\widetilde{\mathcal{O}}\big{(}d\sqrt{{\kappa}/{T}}\big{)},$
where $\rho$ is the context distribution and $\pi^{*}$ is the optimal policy.
@@
We attain the same sub-optimality gap as  Das et al. (2024), but improve the computational efficiency significantly. Our algorithm has an $\mathcal{O}(1)$ time and space complexity per round, while their MLE estimator needs $\mathcal{O}(t\log t)$ time and $\mathcal{O}(t)$ space complexity at iteration $t$.
5.3 Online RLHF with Deployment-Time Adaptation
In this section, we consider the deployment-time adaptation setting, where users provide input contexts in an online manner, and the learner generates responses while simultaneously collecting feedback to improve the model. In this scenario, the learner faces a dual objective: selecting actions that maximize rewards to ensure a positive user experience, while also choosing actions that yield informative feedback to facilitate continual model improvement. To this end, we choose the cumulative regret as the performance measure, defined as
${\operatorname*{Reg}}_{T}=\sum_{t=1}^{T}\left(r\left(x_{t},\pi^{*}(x_{t})\right)-\frac{1}{2}\left(r\left(x_{t},a_{t}\right)+r\left(x_{t},a^{\prime}_{t}\right)\right)\right),$ (8)
where $\pi^{*}$ is the optimal policy. This measure differs from the one introduced by Ji et al. (2024), which evaluates the gap between th e optimal action and a single selected action. Our measure is designed to ensure that both actions are sufficiently effective, thereby delivering a high-quality user experience. While better aligns with deployment needs, it also poses greater challenges. Specifically, to optimize the measure in Ji et al. (2024), the learner can exploit one high-reward action while using the other for uniform exploration. In contrast, our setting imposes a stricter requirement: both actions must be good, thereby limiting the ability to sacrifice one action purely for exploration.
Action selection.  At each iteration, given a new preference data $(x_{t},a_{t},a_{t}^{\prime},y_{t})$, the learner updates the reward model using our one-pass reward modeling method as in Eq. (5). The learner must select queries that are both informative and with high rewards. To address this, we choose the first action $a_{t+1}$ to maximize the estimated reward based on the estimator $\widetilde{\theta}_{t+1}$, i.e.,
$a_{t+1}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\phi(x_{t+1},a)^{\top}\widetilde{\theta}_{t+1}.$ (9)
The second action $a_{t+1}^{\prime}$ aims to maximize the reward and the distance between the two actions, i.e.,
$a_{t+1}^{\prime}=\operatorname*{arg\,max}_{a^{\prime}\in\mathcal{A}}~{}\big{\{}\phi(x_{t+1},a^{\prime})^{\top}\widetilde{\theta}_{t+1}+\widetilde{\beta}_{t+1}\lVert\phi(x_{t+1},a^{\prime})-\phi(x_{t+1},a_{t+1})\rVert_{\mathcal{H}_{t+1}^{-1}}\big{\}}.$ (10)
Theorem 3.
For any $\delta\in(0,1]$, set parameters as in Lemma 1, Algorithm 4 ensures with probability at least $1-\delta$, the regret satisfies
@@ 5.3 Online RLHF with Deployment-Time Adaptation
Our result improves upon Saha et al. (2023) in both computational and statistical efficiency. Statistically, Theorem 3 improves their $\widetilde{\mathcal{O}}\big{(}d{\kappa}\sqrt{{T}}\big{)}$ result by a factor of $\sqrt{\kappa}$. Computationally, our algorithm has an $\mathcal{O}(1)$ time and space complexity per round, while their MLE estimator needs $\mathcal{O}(t\log t)$ time and $\mathcal{O}(t)$ space complexity at iteration $t$ due to optimization over the historical data.
6 Practical Implementation
In this section, we introduce the practical implementation of our proposed algorithm.
6.1 Computation of Inverse Hessian
The OMD update in Eq. (5) requires computation of the inverse of the Hessian matrix. Omitting the projection operation, Eq. (5) can be rewritten as $\widetilde{\theta}_{t+1}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$. Computing the full Hessian inverse ~$\widetilde{\mathcal{H}}_{t}^{-1}$ directly incurs a time complexity of $\mathcal{O}(d^{3})$, which is prohibitive for LLMs.
This cost can be reduced to $\mathcal{O}(d^{2})$ by applying the Sherman-Morrison-Woodbury formula, leveraging the fact that the Hessian is a rank-one update. Specifically, for a matrix of the form $A+\mathbf{x}\mathbf{x}^{\top}$ where $A$ is invertible and $\mathbf{x}$ is a vector, the inverse is given by $(A+\mathbf{x}\mathbf{x}^{\top})^{-1}=A^{-1}-\frac{A^{-1}\mathbf{x}\mathbf{x}^{\top}A^{-1}}{1+\mathbf{x}^{\top}A^{-1}\mathbf{x}}$, requiring only $\mathcal{O}(d^{2})$ time. Nevertheless, even this reduced complexity is costly for large models.
To further reduce the computational burden to $\mathcal{O}(d)$, we employ the Hessian-vector product technique combined with conjugate gradient descent (Boyd and Vandenberghe, 2004). Instead of explicitly computing ~$\widetilde{\mathcal{H}}_{t}^{-1}$, we solve the linear system ~$\widetilde{\mathcal{H}}_{t}v=g_{t}(\widetilde{\theta}_{t})$ iteratively, where $v$ is the solution we seek. The HVP operation ~$\widetilde{\mathcal{H}}_{t}v$ for any vector $v$ can be computed efficiently using twice differentiation: ~$\widetilde{\mathcal{H}}_{t}v=\nabla_{\theta}\big{(}\nabla_{\theta}\mathcal{L}_{t}(\theta)^{\top}v\big{)}\big{|}_{\theta=\widetilde{\theta}_{t}}+\lambda_{t}v$, where $\mathcal{L}_{t}(\theta)$ is the function whose Hessian matrix at $\widetilde{\theta}_{t}$ is ~$\widetilde{\mathcal{H}}_{t}$. By leveraging this efficient computation, the conjugate gradient method iteratively refines the solution. In practice, finding $\mathcal{L}_{t}(\theta)$ is challenging as ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$ involve complex structure. A key insight is ~$\widetilde{\mathcal{H}}_{t}$ can be viewed as the cumulative Hessian of the $\ell_{t}(\theta)$ approximately. This motivates the use of a more tractable surrogate $\ell_{t}(\theta)$ in Eq. (3) while keeping the update. To further approximate the effect of curvature accumulation, we adopt an adaptive damping scheme that approximately preserves historical Hessian information while incorporating new Hessian updates: $\lambda_{t}=\lambda_{0}\cdot\min\{1,f(t/T)\}$, where $f(\cdot)$ is a monotonic increasing function, such as linear functions.
6.2 Computation of Model Uncertainty
In both online RLHF with active data collection and deployment-time adaptation, our algorithm utilizes uncertainty-driven query selection strategies. While quantifying uncertainty using the local norm induced by the Hessian matrix offers strong theoretical guarantees, it is computationally prohibitive in practice. To address this challenge, we adopt a rejection sampling-based approximation, a technique commonly employed for exploration in the literature (Nakano et al., 2021; Gulcehre et al., 2023; Dong et al., 2023, 2024). Specifically, given a prompt, we sample $n$ independent responses by the current model, then use the trained reward model to rank the responses. Then, we use different strategies to select the response for different settings. Specifically, In active data collection, the key insight is to identify and query samples that exhibit the greatest diversity in prompt action features. To this end, we select the response with the highest predicted reward and the response with the lowest predicted reward. In deployment-time adaptation, the core idea is to select the first action to maximize the estimated reward, while the second is chosen to balance high reward with sufficient divergence from the first. Concretely, we select the response with the highest predicted reward and another from the top-$1/q$ percentile of the reward, where $q$ is a hyperparameter.
7 Experiments
@@
In this section, we empirically evaluate the performance of our proposed method.  [^1] Equal contribution. We first describe the experimental setup, and then present the empirical results.
7.1 Experiment setup
@@
In our experiments, we employ the Llama-3-8B-Instruct 222huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and Qwen2.5-7B-Instruct 333huggingface.co/Qwen/Qwen2.5-7B-Instruct as the foundation model for reward model. We extract features $\phi(x,a)$ using the last layer of the model, and the dimension is $d=4096$. We use two datasets for evaluation. The first one is Ultrafeedback-binarized dataset 444huggingface.co/datasets/OpenRLHF/preference_dataset_mixture2_and_safe_pku, a pre-processed version of the original Ultrafeedback dataset (Cui et al., 2023), a widely used benchmark for RLHF. It collects about $64,000$ prompts from diverse resources, including question answering, summarization, and dialogue generation. Each data consists of a context $x$, two responses $a$ and $a^{\prime}$, and a preference label $y$. We also employ a mixed dataset, Mixture2 dataset 444huggingface.co/datasets/OpenRLHF/preference_dataset_mixture2_and_safe_pku, which combines a variety of preference datasets, including HH-RLHF, SHP, UltraFeedback, etc.
@@
11footnotetext: The code is available at https://github.com/ZinYY/Online_RLHF
@@
Figure 2: For Llama-3-8B-Instruct on Ultrafeedback-binarized dataset with passive data collection, we report the comparison of MLE and our OMD-based method for the reward model about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
(a) training loss
(b) training accuracy
(c) evaluation loss
(d) evaluation accuracy
Figure 3: For Qwen2.5-7B-Instruct on Ultrafeedback-binarized dataset with passive data collection, we report the comparison of MLE and our OMD-based method for the reward model about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
(a) training loss
(b) training accuracy
(c) evaluation loss
(d) evaluation accuracy
Figure 4: For Llama-3-8B-Instruct on Mixture2 dataset with passive data collection, we report the comparison of MLE and our OMD-based method for the reward model about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
7.2 Experimental results
In this section, we present the experimental results, including passive data collection, active data collection and deployment-time adaptation.
7.2.1 Passive Data Collection
We randomly sample $T=30,000$ data points from the dataset for training the reward model. We evaluate both the standard MLE-based method, which uses a stochastic gradient descent (SGD) optimizer, and our proposed OMD-based method in terms of reward model loss and accuracy. Figures 23, and 4 present the results for three settings: (i) the Llama-3-8B-Instruct model on the Ultrafeedback-binarized dataset, (ii) the Qwen2.5-7B model on the same dataset, and (iii) the Llama-3-8B-Instruct model on the Mixture2 dataset. We report the results across four key metrics: (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy. Our method demonstrates faster convergence to a lower loss and achieves higher evaluation accuracy compared to the MLE baseline. These results highlight the superior statistical efficiency of the OMD-based approach, which achieves improved performance with fewer training samples.